Write a linear equation through the point (-1,-5) perpendicular to the line y=-3x + 4
Answer by midwood_trail(310)

Write a linear equation through the point (-1,-5) perpendicular to the line
y = -3x + 4
We are looking for another equation.
The quation we are looking for must have a slope that is the negative reciprocal of your given equation.
We are given a point and from the given equation we know the slope is -3.
We use the point-slope formula to find our equation.
(y - y1) = m (x - x1), where m = slope.
y - (-5) = -3(x - (-1))
y + 5 = -3(x + 1)
y + 5 = -3x -3
y = -3x - 3 - 5
y = -3x - 8
The only thing we need to do now is change -3 into 1/3.
The equation that is perpendicular to your given equation is
y = (1/3)(x) - 8